OpenStudy (meilendurcer):

If the lengths of the legs in a right triangle are 7 and 8, what is the approximate length of the hypotenuse?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (sw050399):

Use the formula @audreynicole47 gave you. Pythagorean Theorem! The legs are either a or b (it doesn't which is which) and the hypotenuse is c. So, just plug your numbers in and simplify! To set it up: 7^2 + 8^2 = c^2 Solve for c.

OpenStudy (meilendurcer):

113?

OpenStudy (anonymous):

Yes. That is c^2. Then you square both sides and solve.

OpenStudy (mertsj):

Now just take the square root of that number because c^2=113, so c + ??

OpenStudy (anonymous):

It should be about 10.6